OpenStudy (afloridagirl):

A firecracker shoots up from a hill 140 feet high, with an initial speed of 100 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Five seconds Seven seconds Nine seconds 11 seconds

jimthompson5910 (jim_thompson5910):

`initial speed of 100 feet per second` means v = 100

jimthompson5910 (jim_thompson5910):

The initial height is 140 ft because `A firecracker shoots up from a hill 140 feet high` so s = 140

jimthompson5910 (jim_thompson5910):

\[\Large H(t) = -16t^2+vt+s\] turns into \[\Large H(t) = -16t^2+100t+140\] making sense so far?

OpenStudy (afloridagirl):

so in so..

OpenStudy (afloridagirl):

i think it would be 7 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Did you use the quadratic formula to solve?

OpenStudy (afloridagirl):

so in so...

OpenStudy (afloridagirl):

i thought it would be 7 or 9

jimthompson5910 (jim_thompson5910):

did you get a decimal result?

jimthompson5910 (jim_thompson5910):

I'm curious to see what you got before you rounded

OpenStudy (afloridagirl):

tbh it was a guess..

jimthompson5910 (jim_thompson5910):

Looking at \[\Large H(t) = -16t^2+100t+140\] do you agree that a = -16 b = 100 c = 140 ??

OpenStudy (afloridagirl):

whats a,b,c

OpenStudy (afloridagirl):

like howd you get that?

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{-16}}t^2+{\color{green}{100}}t+{\color{blue}{140}}\] is in the form \[\Large {\color{red}{a}}t^2+{\color{green}{b}}t+{\color{blue}{c}}\] where \[\Large \color{red}{a = -16}\] \[\Large \color{green}{b = 100}\] \[\Large \color{blue}{c = 140}\]

OpenStudy (afloridagirl):

oh

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (afloridagirl):

yes

jimthompson5910 (jim_thompson5910):

so you'll plug a = -16 b = 100 c = 140 into the quadratic formula \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (afloridagirl):

okay

jimthompson5910 (jim_thompson5910):

tell me what two results you get. You should get decimal values

OpenStudy (afloridagirl):

ik how to plug them in but i dont know how to solve... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You have this so far? \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-100\pm\sqrt{100^2-4*(-16)*(140)}}{2*(-16)}\]

OpenStudy (afloridagirl):

so we solve?

jimthompson5910 (jim_thompson5910):

what is 100^2 equal to?

OpenStudy (afloridagirl):

10000

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how about 4*(-16)*(140) ?

OpenStudy (afloridagirl):

-8960?

jimthompson5910 (jim_thompson5910):

So \[\Large x = \frac{-100\pm\sqrt{100^2-4*(-16)*(140)}}{2*(-16)}\] turns into \[\Large x = \frac{-100\pm\sqrt{10,000-(-8,960)}}{2*(-16)}\]

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-100\pm\sqrt{10,000-(-8,960)}}{2*(-16)}\] turns into \[\Large x = \frac{-100\pm\sqrt{10,000+8,960}}{2*(-16)}\]

OpenStudy (afloridagirl):

and the bottom is -32?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (afloridagirl):

okay so the top is -1140 and the bottom is -32?

jimthompson5910 (jim_thompson5910):

10,000+8,960 = ???

OpenStudy (afloridagirl):

18960

jimthompson5910 (jim_thompson5910):

take the square root of that to get what?

OpenStudy (afloridagirl):

??

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{18,960} = ???\]

jimthompson5910 (jim_thompson5910):

type `sqrt(18960)` into a calculator http://web2.0calc.com/

OpenStudy (afloridagirl):

okay

jimthompson5910 (jim_thompson5910):

tell me what you get for the result of sqrt(18960)

OpenStudy (afloridagirl):

137.695315824468045

jimthompson5910 (jim_thompson5910):

good

OpenStudy (afloridagirl):

im sorry this is taking forever

jimthompson5910 (jim_thompson5910):

so we have this now \[\Large x \approx \frac{-100\pm137.695315824468045}{-32}\]

jimthompson5910 (jim_thompson5910):

that's ok. It's good not to rush

OpenStudy (afloridagirl):

so we divide -32

jimthompson5910 (jim_thompson5910):

break up the plus/minus to get these two equations \[\Large x \approx \frac{-100+137.695315824468045}{-32}\] OR \[\Large x \approx \frac{-100-137.695315824468045}{-32}\]

jimthompson5910 (jim_thompson5910):

Simplify the top first, then divide by -32

OpenStudy (afloridagirl):

okay i got 4.3..

jimthompson5910 (jim_thompson5910):

let's focus on the "plus" first \[\Large x \approx \frac{-100+137.695315824468045}{-32}\] \[\Large x \approx \frac{37.695315824468}{-32}\] \[\Large x \approx -1.17797861951462\] \[\Large x \approx -1.178\]

jimthompson5910 (jim_thompson5910):

agreed? or no?

OpenStudy (afloridagirl):

yes

jimthompson5910 (jim_thompson5910):

ok let's focus on the "minus" now \[\Large x \approx \frac{-100-137.695315824468045}{-32}\] \[\Large x \approx \frac{-237.695315824468}{-32}\] \[\Large x \approx 7.42797861951462\] \[\Large x \approx 7.428\]

OpenStudy (afloridagirl):

so seven?

jimthompson5910 (jim_thompson5910):

So the two solutions to \[\Large -16t^2+100t+140=0\] are \[\Large t \approx -1.178 \ \ \text{ or } \ \ t \approx 7.428\]

jimthompson5910 (jim_thompson5910):

the negative time value makes no sense, which means we ignore it

jimthompson5910 (jim_thompson5910):

yes approx 7 seconds

OpenStudy (afloridagirl):

okay thank you im sorry that i took so long

jimthompson5910 (jim_thompson5910):

that's ok